OpenStudy (anonymous):

19) The price of a calculator has decreased from $200 to $150. What is the percentage of decrease? A) 3% B) 25% Eliminate C) 33% D) 50%

OpenStudy (blake57roger):

d

OpenStudy (anonymous):

B

OpenStudy (shamim):

200-150=?

OpenStudy (blake57roger):

B

OpenStudy (anonymous):

200 * .25 = 50. 200 - 50 = 150

OpenStudy (trojanpoem):

50 * 100 / 200 = 25 percent

OpenStudy (blake57roger):

Its not d Whats 250 divided by 4

OpenStudy (blake57roger):

25%

OpenStudy (blake57roger):

metal

OpenStudy (anonymous):

ok

OpenStudy (shamim):

Why u r telling result????!!!!!!!